For the confidence interval below, what is the sample mean, ?
18 < u < 24
1
Expert's answer
2014-07-21T11:57:09-0400
The sample mean is the average of endpoints of confidence interval, i.e. the sample mean is (18+24)/2=42/2=21.
Answer: 21.